# A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations

### v1

@@ Abstract
Accurate maps of atrial electrical activation are essential for personalised treatment of arrhythmias, yet biophysically detailed simulations remain computationally intensive for real-time clinical use or population-scale analyses. Here we introduce a geometry-independent operator-learning framework that predicts local activation time (LAT) fields across diverse left atrial anatomies with near-instantaneous inference. We generated a dataset of 308,700 simulations using a GPU-accelerated electrophysiology solver, systematically varying multiple pacing sites and physiologically varied conduction properties across 147 patient-specific geometries derived from two independent clinical cohorts. All anatomical and functional data are expressed in a Universal Atrium Coordinate system, providing a consistent representation that decouples electrophysiological patterns from mesh topology. Within this coordinate space, we designed a neural operator with a vision-transformer backbone to learn the mapping from structural and electrophysiological inputs to LAT fields. With a mean prediction error of 5.1 ms over a 455 ms maximum simulation time, the model outperforms established operator-learning approaches and performs inference in 0.12 ms per sample. Our framework establishes a general strategy for learning domain-invariant biophysical mappings across variable anatomical domains and enables integration of computational electrophysiology into real-time and large-scale clinical workflows.
@@ 1 Introduction
Atrial fibrillation (AF) is the most common sustained cardiac arrhythmia, affecting around 46.3 million individuals worldwide [kornej2020epidemiology], and represents a major cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes]. Computational simulations have emerged as powerful tools in AF research and clinical trials, offering insights into arrhythmia mechanisms, enabling the prediction of ablation outcomes, and personalised therapeutic strategies [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects]. Many of these applications rely on calibrating patient-specific biophysical models to local activation time (LAT) maps, which capture the spatiotemporal propagation of electrical wavefronts across the myocardium. This calibration typically involves solving large ensembles of high-fidelity finite element method (FEM) simulations that numerically solve the governing partial differential equations (PDEs) of cardiac electrophysiology on anatomically accurate patient-specific geometries [goktepe2009computational, rocha2010macro].
Although highly accurate, FEM workflows are computationally expensive and often take hours to days to solve, making them impractical for intraoperative use. Consequently, patient-specific models are usually prepared in advance from pre-procedure MRI or CT scans that define anatomy and tissue properties. In many clinical scenarios, however, such imaging is unavailable, and the only anatomical and electrophysiological data come from intraoperative electro-anatomical mapping (EAM). While EAM-derived LATs can be used for retrospective calibration [jaffery2024review], the turnaround time of FEM-based methods prevents their use for real-time decision-making during ablation.
Recent advances in machine learning have opened new opportunities to accelerate the numerical solution of cardiac electrophysiology models. Physics-informed neural networks (PINNs) have been proposed as a physics-constrained alternative to purely data-driven surrogates, embedding the governing PDE residuals directly into the loss function [pinn]. They are typically trained for a single PDE configuration, defined by a fixed geometry, parameter set, and boundary conditions [kovachki2023neural]. However, they do not naturally generalise to new configurations without retraining: if the domain shape, material parameters, or boundary conditions change, the PINN must be re-trained from scratch or at least fine-tuned.
Neural operators directly address this limitation by learning a mapping between infinite-dimensional function spaces rather than a single instance of a PDE [kovachki2023neural]. Foundational neural operators such as the Deep Operator Network (DeepONet) [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on canonical problems defined over structured domains. By learning this operator mapping, they can efficiently predict solution fields for new inputs, even for unseen geometries or different discretisation resolutions, without the need for time-consuming retraining. Applying neural operators to complex real-world problems, such as predicting cardiac activation maps, presents significant challenges. First, these aforementioned neural operator frameworks are built to handle data on structured, canonical domains, making them ill-suited for diverse and irregular geometries [geofno, gino]. Second, training a reliable neural operator is contingent on the availability of large, high-quality datasets of diverse simulation results [ahmad2024diffeomorphic], and generating such datasets is a major challenge [trayanova2021machine].
Several approaches have been proposed to address the challenge of anatomical variability. Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable] uses diffeomorphic transformations to map individual geometries to a shared, standardised space. This allows a neural operator to learn latent solutions on the reference domain, with the original shape encoded as an auxiliary input. While effective for relatively simple geometries, such as the semi-spherical left ventricle (LV), this strategy faces substantial difficulties when extended to anatomies with more intricate topologies, such as the multi-holed structure of the left atrium, where constructing topologically consistent reference domains is non-trivial. Similarly, methods like GeoFNO require a diffeomorphic mapping to simple domains (e.g., cubes or tori) where the Fourier transform can be applied [geofno]; however, the topologically complex 3D geometries of the human left atrium, with multiple openings, make such mappings impossible [loeffler2024graph]. Graph-based methods like the Graph Fourier Neural Kernel (G-FuNK) [loeffler2024graph] learn directly on mesh graphs, avoiding domain mapping. However, their spectral bases are sensitive to mesh quality and topology, reducing robustness and generalisability in heterogeneous clinical datasets.
Here, we introduce a method that overcomes the dependency of neural operators on structured domains for electrophysiology simulations. Our approach projects inputs and outputs onto a common, anatomically standardised reference using the Universal Atrial Coordinates (UAC) system [uac]. The UAC defines a mesh-independent coordinate space for the left atrium (LA), preserving physiologically relevant spatial relationships while unifying geometry and topology across patients. This representation enables neural operators such as DeepONet and FNO to consistently interpret spatial features, free from irregular mesh geometries or topological inconsistencies.
To support large-scale training, we developed a GPU-accelerated FEM solver, drastically reducing the time required to compute LAT maps and enabling the generation of a high-fidelity, comprehensive dataset. Using this dataset, we train a neural operator to approximate the underlying biophysical mapping, achieving near real-time inference without compromising physiological accuracy. Our model, an encoder-decoder architecture with a Vision Transformer (ViT) [vit] backbone, predicts full-field LAT maps from patient-specific anatomical features, conductivity, and pacing sites. It achieves a mean absolute error of $5.1\pm 0.2$ ms relative to FEM simulations, exhibiting high pointwise accuracy, spatial smoothness, and robust generalisation across independent clinical cohorts, overcoming the domain shift inherent in multi-centre medical data.
By integrating GPU-accelerated simulation and neural operator training into a unified framework, our approach addresses both the computational bottleneck of large-scale electrophysiology simulations and the geometric variability of patient-specific atrial anatomies. It enables accurate prediction of full-field LAT maps in just 0.12 ms per sample on a GPU, offering a pathway toward rapid, patient-specific, physiology-informed decision support for AF interventions.
@@ 2.1 Large-scale EP simulations on GPUs
To assess potential cohort-level differences, we performed dimensionality reduction on the full set of UAC-projected LAT maps. High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform]. The 2D embedding (Fig. 1b) reveals a clear separation between the two datasets. Within each cohort, seven well-defined clusters emerge, each corresponding to one of the pacing sites. Clusters associated with the same pacing location in Datasets A and B do not overlap but instead form displaced pairs in the latent space. This consistent displacement provides visual evidence of a domain shift between the cohorts, arising not from physiological differences, but from systematic variations in anatomical distributions, acquisition protocols, or preprocessing pipelines. This domain shift has a measurable impact on cross-domain generalisation performance.
@@ 2.2 Input feature sensitivity analysis
We evaluated the contribution of each anatomical input modality by removing spatial coordinates, fibre orientation, and surface area features in turn, and retraining the model under each condition. Performance was assessed using two complementary metrics: Mean Absolute Error (MAE), which quantifies the numerical accuracy (error in milliseconds), and the Structural Similarity Index Measure (SSIM) [wang2004image], which measures the spatial fidelity and structural quality of the predicted activation maps. All analyses in this section were performed using 5-fold cross-validation. Unless stated otherwise, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
The results in Fig. 2a show that incorporating all three feature types achieves accurate and spatially smooth LAT predictions (MAE $6.09\pm 0.26$ ms; SSIM $0.963\pm 0.001$). Removing the surface area feature results in a pronounced degradation in performance (MAE $9.88\pm 0.81$ ms), indicating that global anatomical scaling is an essential contextual signal for correct wavefront organisation. Similarly, excluding spatial coordinates reduces both numerical accuracy and structural similarity (MAE $9.04\pm 0.44$ ms; SSIM $0.948\pm 0.002$), demonstrating that explicit localisation in the UAC domain is critical for resolving conduction pathways.
In contrast, removing fibre orientation led to improved performance across both metrics (MAE $5.15\pm 0.25$ ms; SSIM $0.978\pm 0.002$). This suggests that the fibre orientation information, as currently represented, may introduce noise or redundancy relative to the geometric cues already encoded in the coordinate and surface area channels. We discuss this observation further in the discussion section. Accordingly, fibre orientation features were excluded from all subsequent experiments.
@@ 2.3 Evaluating the effect of regularisation
To assess the impact of regularisations on LAT map reconstruction, we evaluated the contribution of two common spatial regularisation terms, total variation and second-order Laplacian loss, when added individually or in combination. These regularizers aim to encourage spatial smoothness and suppress high-frequency artefacts in the predicted LAT maps.
To evaluate the impact of different loss formulations on model performance, we conducted systematic experiments by varying the regularisation coefficient $\lambda\in\{5\times 10^{-2},3\times 10^{-2},1\times 10^{-2},5\times 10^{-3},1\times 10^{-3}\}$ in the loss function.
For each fold in 5-fold cross-validation, the model achieving the best performance on the validation set was retained, and the corresponding metrics were recorded. The performance for a given regularisation weight $\lambda$ was derived by averaging these metrics across the five folds. To probe the effect of each regularisation configuration, such as TV alone or the combination of TV and Laplacian, we further averaged the fold-averaged metrics across all tested $\lambda$ values (Fig. 2b).
Among all evaluated configurations, the $H^{1}$ loss combined with both TV and Laplacian regularisations achieved the best overall performance, with a mean MAE of $5.16$ ms and a mean SSIM of $0.978$. This demonstrates that combining both regularisation terms effectively enhances structural fidelity while preserving numerical accuracy. The addition of TV or Laplacian individually yielded smaller improvements, suggesting that their combined effect is necessary to fully capture the relevant spatial structures.
@@ 2.4 Benchmarking model performance
We benchmarked a range of neural operator frameworks and deep-learning architectures capable of modelling cardiac activation patterns from multimodal inputs. For each model, we selected three primary hyperparameters governing capacity and constructed six configurations by varying these across representative ranges, ensuring that the comparison is both systematic and approximately capacity-matched across architectures (Fig. 2c). Supplementary Section 2 provides further details on architecture design and training procedures.
We first evaluated operator learning approaches. While the Fourier Neural Operator (FNO) [fno] is theoretically well-suited for learning global mappings in function space, its empirical performance was moderate (MAE $10.14\pm 0.07$ and SSIM $0.938\pm 0.005$), indicating potential challenges in handling heterogeneous and noisy physiological data. The Wavelet Neural Operator (WNO) [wno] exhibited the weakest performance among evaluated models (MAE $10.08\pm 0.28$ and SSIM $0.938\pm 0.007$, likely reflecting sensitivity to measurement noise and anatomical variability inherent in real-world datasets. The DeepONet framework, using our proposed ViT-based architecture for both branch and trunk networks, achieved strong performance (MAE $7.12\pm 0.37$ ms; SSIM $0.963\pm 0.005$).
@@
Among convolutional baselines, the U-Net architecture [unet] outperformed the simpler ResNet model [resnet]. The U-Net’s multi-scale encoder-decoder structure, with its strong inductive bias for spatial localisation, yielded a balanced performance of MAE $8.48\pm 0.28$ ms and SSIM $0.965\pm 0.004$. However, these convolutional models were fundamentally limited in their ability to capture global pacing effects and long-range spatial dependencies
Our model surpassed all baselines, achieving the best results with an MAE of $5.15\pm 0.25\text{ ms}$ and SSIM of $0.978\pm 0.002$. This confirms the model’s superior capability to synthesise heterogeneous inputs into spatially coherent, smooth, and highly accurate LAT map predictions.
While MAE and SSIM provide global measures of model performance, they do not reveal where errors concentrate spatially. We therefore visualised the prediction error distributions across the UAC domain. Fig. 2d displays the spatial distribution of prediction error on LAT maps for six test cases, each averaged over seven pacing locations and 300 combinations of conductivity velocities. Across all six test cases, a visually consistent and highly coherent source of localised prediction error is observed near the horizontal central line of the UAC map. This localised band of elevated error is the most prominent spatial feature in the error distribution plots. Beyond this central band, the prediction errors generally remain low and diffuse across the rest of the UAC domain, indicating high model fidelity in most tissue regions.
@@
Beyond predictive accuracy, we further assessed the computational efficiency of our model to evaluate its suitability for real-time applications. Inference on batches of 1000 samples was performed on an NVIDIA RTX 5090 GPU in evaluation mode with CUDA synchronisation to ensure precise wall-clock timing. The model achieved an average inference time of $122.23$ ms per batch, or approximately $0.12$ ms per sample. This represents a five to six orders of magnitude acceleration compared with conventional finite-element electrophysiology solvers, which require approximately one minute per simulation with our solver in this study. Such efficiency enables near real-time prediction of LAT maps across diverse atrial anatomies and pacing configurations, highlighting the potential of our model for integration into time-critical clinical workflows and large-scale population studies.
@@ 2.5 Model generalisation across pacing sites
Understanding whether a single model can generalise across diverse pacing conditions is critical for building scalable and clinically useful predictors of atrial activation, as it allows us to determine whether it can match or exceed the performance of individually trained models. Success in this setting would indicate that the model captures shared electrophysiological structure, enabling robust predictions without prior knowledge of stimulation location.
@@ 2.6 The challenge of latent data heterogeneity
Each evaluation was performed on models trained with 18 combinations of hyperparameters to ensure statistical significance, with results highlighting consistent trends (Fig. 2f; full results in Supplementary Section 3). The strong internal performance confirms that the UAC projection effectively separates anatomical representation from patient-specific shape. In contrast, the pronounced drop in cross-domain performance provides clear evidence of latent bias. The model relies on image-specific cues that mislead predictions when applied to a new domain. This finding highlights a second-order challenge in translational modelling. Models are highly sensitive to systematic variations in data representation and annotation rather than true anatomical differences.
Integrating heterogeneous datasets during training mitigates these domain-specific biases. Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines. Across both datasets, MAE decreased and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation. For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.

### v2

@@ Abstract
Learning biophysically accurate solution operators for cardiac electrophysiology is fundamentally challenged by geometric variability across patient-specific heart anatomies. Most existing neural operator approaches are limited to structured or weakly deformed domains, restricting their applicability to realistic atrial and ventricular geometries. Here, we introduce a unified operator-learning framework that projects inputs and outputs onto a standardised anatomical coordinate system, decoupling electrophysiological dynamics from mesh topology. This formulation enables geometry-independent learning while preserving physiologically meaningful spatial organisation, and allows predictions to be interpolated back onto patient-specific geometries for anatomical interpretation.
To support large-scale training within the framework, we develop a GPU-accelerated electrophysiology solver and generate over 300,000 high-fidelity simulations across diverse patient-specific left atrial geometries with varied pacing and conduction properties. Within this anatomical coordinate domain, we design a neural operator to predict full-field local activation time maps, achieving a mean absolute error of 5.1 ms and an inference time of 0.12 ms per sample, outperforming existing operator learning and convolutional baselines. We further validate the framework on ventricular geometries, demonstrating robust generalisation beyond the atrial setting. Together, this framework establishes a scalable foundation for fast, geometry-invariant cardiac electrophysiology modelling, with potential relevance for real-time and population-scale clinical workflows.
@@ 1 Introduction
Computational models of cardiac electrophysiology provide a principled framework for studying electrical wave propagation, investigating arrhythmia mechanisms, and informing patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized]. A central application of these models is the analysis of local activation time (LAT) maps, which characterise the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac]. Generating and analysing such LAT maps typically requires large ensembles of high-fidelity finite element method (FEM) simulations on anatomically detailed, patient-specific geometries, involving the numerical solution of coupled partial differential equations governing cardiac electrophysiology [niederer2011verification, opencarp].
Atrial fibrillation (AF) is a primary clinical setting in which such patient-specific LAT analyses are routinely required [sahli2018interpreting]. As the most common sustained cardiac arrhythmia, AF affects approximately 46.3 million individuals worldwide [kornej2020epidemiology] and is a leading cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes]. Computational modelling has therefore become an important tool in AF research and clinical studies, supporting mechanistic investigation of arrhythmia drivers, evaluation of ablation strategies, and development of personalised treatment planning [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects]. However, the substantial computational cost of FEM-based workflows, often requiring hours to days per simulation, limits their applicability in settings that demand rapid or real-time decision-making [niederer2011simulating, boyle2019computationally].
Recent advances in machine learning have opened new avenues for accelerating cardiac electrophysiology simulations. Neural operators, in particular, aim to learn mappings between infinite-dimensional function spaces [kovachki2023neural], enabling solution fields to be predicted for unseen inputs, discretisations, or geometries without retraining. Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on parametric partial differential equation problems. Despite this promise, most existing neural operator frameworks are developed for structured or weakly deformed computational domains, which limits their applicability to the highly irregular, patient-specific geometries encountered in cardiac modelling [geofno, gino]. In addition, training robust and generalisable operators requires large, high-quality simulation datasets, which remain expensive to generate using conventional electrophysiology solvers [ahmad2024diffeomorphic, trayanova2021machine].
Several strategies have been proposed to address anatomical variability in operator learning. Diffeomorphic mapping approaches [zhao2025diffeomorphism], including Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable], map individual geometries to a shared reference domain, enabling learning in a common representation while encoding geometry as auxiliary information. While effective for relatively simple geometries, such as the semi-spherical left ventricle, these approaches become challenging for anatomies with complex topology, such as the multi-holed left atrium, where constructing globally consistent mappings is non-trivial. Related methods, such as Geo-FNO [geofno], deform irregular geometries to structured domains to enable spectral learning, but the accuracy and stability of such mappings may deteriorate for highly intricate anatomical structures.
Here, we propose a framework that reformulates neural operator learning in a coordinate domain intrinsic to cardiac anatomy rather than in the space of mesh discretisations. By expressing model inputs and outputs in a shared anatomical parameterisation, electrophysiological patterns can be represented consistently across patients despite substantial geometric and topological variability. Predictions obtained in this coordinate space can subsequently be mapped back onto the original patient-specific geometry, enabling anatomically faithful reconstruction and direct clinical interpretation. In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] systems to define such intrinsic representations for atrial and ventricular geometries, respectively. These coordinate systems preserve physiologically meaningful spatial organisation while providing a unified domain across diverse anatomies, enabling neural operators to learn spatially coherent solution mappings without sensitivity to irregular mesh connectivity or topological inconsistencies.
To support operator training within this framework, we further develop a GPU-accelerated electrophysiology simulation pipeline that enables efficient generation of high-fidelity LAT maps at scale across diverse patient-specific geometries. By integrating high-performance simulation with neural operator learning in anatomical coordinate space, our approach addresses both the computational cost of cardiac electrophysiology modelling and the geometric variability inherent in patient-specific anatomies, providing a scalable pathway toward rapid, physiology-informed inference in atrial fibrillation.
@@ 2.1 Large-scale EP simulations on GPUs
To assess potential cohort-level differences, we performed dimensionality reduction on the full set of UAC-projected LAT maps. High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform]. The 2D embedding (Fig. 1b) reveals a clear separation between the two datasets. Within each cohort, seven well-defined clusters emerge, each corresponding to one of the pacing sites. Clusters associated with the same pacing location in Datasets A and B do not overlap but instead form displaced pairs in the latent space. This consistent displacement provides visual evidence of a domain shift between the cohorts, arising not from physiological differences, but from systematic variations in anatomical distributions, acquisition protocols, or preprocessing pipelines. This domain shift has a measurable impact on cross-domain generalisation performance. The methodology used to assess this effect is detailed in Methods.
@@ 2.2 Input feature sensitivity analysis
We quantified the contribution of individual anatomical input modalities by systematically removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration. Model performance was evaluated using mean absolute error (MAE), which captures numerical accuracy in milliseconds, and the structural similarity index measure (SSIM) [wang2004image], which assesses spatial fidelity of the predicted activation maps. All results are reported using five-fold cross-validation. Unless otherwise stated, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
As shown in Fig. 2a, incorporating all anatomical features yields accurate and spatially coherent predictions of local activation time. Removing the surface area feature leads to a substantial degradation in performance, highlighting the importance of global anatomical scaling for correctly organising activation wavefronts. Similarly, excluding spatial coordinates significantly reduces both numerical accuracy and structural similarity, demonstrating that explicit localisation in the UAC domain is essential for resolving conduction pathways.
In contrast, removing fibre orientation features unexpectedly improves predictive performance across both metrics. This observation indicates that, in the present formulation, fibre orientation does not provide an additional informative signal beyond that already captured by geometric coordinates and global anatomical features. Based on this result, fibre orientation inputs were excluded from all subsequent experiments. Explanations for this effect are discussed further below.
@@ 2.3 Evaluating the effect of regularisation
We investigated the impact of spatial regularisation on LAT map reconstruction by augmenting the $H^{1}$ loss with two commonly used smoothness terms: total variation (TV) and a second-order Laplacian penalty, applied individually or in combination. These regularisation terms are designed to suppress high-frequency artefacts while preserving physiologically meaningful spatial structure.
Model performance was evaluated across a range of regularisation strengths ($\lambda\in\{5\times 10^{-2},3\times 10^{-2},1\times 10^{-2},5\times 10^{-3},1\times 10^{-3}\}$) using five-fold cross-validation. For each configuration, performance metrics were averaged across folds and regularisation weights to isolate the effect of the regularisation formulation (Fig. 2b).
Among all tested configurations, combining the $H^{1}$ loss with both TV and Laplacian regularisation consistently yielded the best performance, achieving the lowest mean MAE and highest SSIM. This result indicates that jointly enforcing first- and second-order smoothness constraints enhances structural fidelity without sacrificing numerical accuracy. In contrast, applying either TV or Laplacian regularisation alone resulted in smaller and less consistent improvements, suggesting that their combined effect is necessary to adequately capture the spatial organisation of activation patterns.
@@ 2.4 Benchmarking model performance
We benchmarked our approach against a range of neural operators and convolutional architectures commonly used for modelling cardiac activation from multimodal inputs. To ensure a fair comparison, models were evaluated across approximately capacity-matched configurations spanning representative hyperparameter ranges (Fig. 2c). Full architectural details and exhaustive results are provided in Supplementary Section 2 and Supplementary Table 1.
Among operator learning methods, the Fourier Neural Operator (FNO) and Wavelet Neural Operator (WNO) exhibited limited empirical performance in this setting, indicating challenges in capturing heterogeneous and noisy electrophysiological patterns on complex anatomical domains. DeepONet, implemented using our ViT-based backbone for both branch and trunk networks, achieved substantially improved accuracy, reflecting its stronger capacity to model global functional relationships.
@@
Convolutional baselines showed a similar trend. The U-Net architecture outperformed the simpler ResNet model, consistent with its multi-scale encoder–decoder structure and strong inductive bias for spatial localisation. Nevertheless, convolutional models remained limited in their ability to capture global pacing effects and long-range spatial dependencies.
Our model outperformed all baselines, achieving the highest numerical accuracy and structural fidelity across all evaluated metrics. This result demonstrates its superior ability to integrate heterogeneous anatomical and physiological inputs into spatially coherent and physiologically consistent LAT predictions.
To assess where residual errors concentrate spatially, we visualised prediction error distributions across the UAC domain (Fig. 2d). Across representative test cases, errors are predominantly localised to a narrow horizontal band corresponding to the UAC seam, while remaining low and diffuse elsewhere, indicating high fidelity across most anatomical regions.
@@
Beyond accuracy, we evaluated computational efficiency to assess suitability for real-time applications. On a modern NVIDIA GPU, the model achieved an average inference time of approximately $0.12$ ms per sample. This represents orders of magnitude acceleration relative to conventional finite-element electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
@@ 2.5 Model generalisation across pacing sites
Generalisation across diverse pacing conditions is critical for scalable and clinically useful prediction of atrial activation, as it determines whether a single model can match or exceed the performance of models trained for individual pacing sites. A model that performs well across pacing sites indicates that it has learned shared electrophysiological structure rather than site-specific patterns.
@@ 2.6 The challenge of latent data heterogeneity
Each evaluation was performed on models trained with 18 combinations of hyperparameters to ensure statistical significance, with results highlighting consistent trends (Fig. 2f; full results are provided in Supplementary Tables 2 and 3). The strong internal performance confirms that the UAC projection effectively separates anatomical representation from patient-specific shape. In contrast, the pronounced drop in cross-domain performance provides clear evidence of latent bias. The model relies on image-specific cues that mislead predictions when applied to a new domain. This finding highlights a second-order challenge in translational modelling. Models are highly sensitive to systematic variations in data representation and annotation rather than true anatomical differences.
Integrating heterogeneous datasets during training mitigates these domain-specific biases. Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines. Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation. For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
2.7 Interpolation to patient-specific geometry
To assess anatomical prediction fidelity after interpolation, we evaluated model performance on both LA surface meshes and left ventricular (LV) volume meshes (Fig. 3). Predictions produced in the intrinsic anatomical coordinate domain were interpolated back onto the original patient-specific mesh geometry, enabling anatomically faithful reconstruction in both settings.
Figure 3: Interpolation of neural operator predictions to patient-specific atrial and ventricular mesh geometries. a, Left: local activation time (LAT) predictions interpolated from the coordinate domain back onto representative left atrial (LA) surface meshes. Regions where the absolute prediction error exceeds 15 ms are highlighted. Right: distribution of mean absolute error across pacing locations evaluated on 100 LA test cases. b, Corresponding results for left ventricular (LV) meshes, showing interpolated LAT predictions, regions of large absolute error, and error distributions across pacing locations.
Model performance was evaluated using a previously proposed benchmark task in which LAT is predicted from one of seven pacing locations alone [yin2024scalable]. Our experiments were conducted on Cohort A using 100 left atrial geometries, of which 20 were held out for testing. In addition, 1,006 left ventricular (LV) meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation. Model inputs and outputs were represented on a $100\times 100$ grid for LAs and a $50\times 50\times 50$ volumetric grid for LVs.
On test data, the model achieves mean absolute errors of 5–6 ms for LA geometries, relative to maximum activation times of up to 285 ms. For LV geometries, errors are consistently lower, typically ranging between 2 and 4 ms for maximum activation times of up to 350 ms. These ventricular error levels improve upon those reported in prior geometry-independent operator learning approaches, where mean absolute errors of approximately 5–6 ms were observed for LV meshes [yin2024scalable]. In both atrial and ventricular settings, errors remain consistent across pacing locations, indicating stable performance independent of stimulation site. Spatially, elevated errors in the LA are primarily localised to the atrial appendage and pulmonary vein openings, regions characterised by complex geometry and boundary effects. For the LV, larger errors are observed near the basal openings, corresponding to truncation at the valve plane, and along the interface with the right ventricle, reflecting geometric discontinuities introduced during extraction from bi-ventricular meshes.